OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1

OpenStudy (anonymous):

y=3 x=1

OpenStudy (anonymous):

Because: (-2)(-2x+y=1) = 4x-2y=-2. Use this equation and the second equation to cancel out the 4x's. Then you get -y=-3, and then y=3. Then you plug y in to solve for x, and get -2x + (3) = 1. Then -2x=-2, then x=1